Sally travels by car from one city to another. She drives for 22.0 min at 69.0 km/h, 49.0 min at 44.0 km/h, and 30.0 min at 36.0 km/h, and she spends 11.0 min eating lunch and buying gas.
1
Expert's answer
2018-09-20T10:55:09-0400
(a) The average speed is
v=d/t=(v_1 t_1+v_2 t_2+v_3 t_3+v_4 t_4)/(t_1+t_2+t_3+t_4 )
v_4=0. So,
v=(v_1 t_1+v_2 t_2+v_3 t_3)/(t_1+t_2+t_3+t_4 )
v=((69)(22/60)+(44)(49/60)+(36)(30/60))/((22/60)+(49/60)+(30/60)+11/60)=42.4 km/h.

(b) The total distance traveled:
d=v_1 t_1+v_2 t_2+v_3 t_3
d=(69)(22/60)+(44)(49/60)+(36)(30/60)=79.2 km.